THE GLYCINE-WATER DIMER

\underline{JOSÉ L. ALONSO}, EMILIO J. COCINERO, ALBERTO LESARRI and JUAN C. LÓPEZ, Grupo de Espectroscopía Molecular (GEM), Departamento de Química Física y Química Inorgánica, Facultad de Ciencias, Universidad de Valladolid, E-47005 Valladolid, Spain.

We report the first observation of the 1:1 complex of glycine-water. A solid rod of glycine was impinged by the green line of a Nd:YAG laser and the vaporized molecules were seeded into a gas pulse of Ne saturated with water, forming a supersonic jet which was probed by FT-microwave spectroscopy.^{a} The water molecule has been reliably located in the complex from the analysis of several isotopomers. Glycine is in its most stable form, with a cis-carboxylic group and a NH⋯O=C bifurcated intramolecular hydrogen bond. The complex is stabilized by two intermolecular hydrogen bonds formed between the carbonyl group and one of the hydrogen atoms of water (O_{w}-H⋯O=C) and between the hydroxyl group and the electron lone pair at the oxygen atom of water (O_{w} ⋯H-O-C)).

aA. Lesarri, S. Mata, J. C. López, and J. L. Alonso; Rev. Sci. Instrum., 74, 4799 (2003)